KyledaGreat:

Find a formula for the inverse of the following function, if possible. W(x)=4x^1/5+3 Selecting a radio button will replace the entered answer value(s) with the radio button value. If the radio button is not selected, the entered answer is used. W^-1(x) = ___________________ does not have an inverse function

KyledaGreat:

@vocaloid

minustempo:

You find the inverse by essentially just switching the x and y places

KyledaGreat:

\[W(x) = 4x^{\frac{ 1 }{ 5}} + 3\] that's what it look like

KyledaGreat:

W(x) 4y^1/5 + 3 you mean ?

KyledaGreat:

w = 4/y^4/5 + 3/y

minustempo:

Sorry you can substitute y = W(x) so y = 4x^{1/5} + 3 so x = 4y^{1/5} + 3

minustempo:

so 4y^(1/5) = x-3 y^(1/5) = (x-3)/4

KyledaGreat:

W^-1(x) = y^(1/5) = (x-3)/4 is the answer ?

minustempo:

no y is the inverse y^(1/5) = (x-3)/4 y = ((x-3)/4)^5 so ((x-3)/4)^5 is the inverse

KyledaGreat:

it wasn't right

minustempo:

what's the answer?

KyledaGreat:

it don't say

minustempo:

@sammixboo I don't think I'm wrong but would you help